Find side a in ∆ABC, side b = 12 m, side c = 20 m and angle B = 45˚.


1
Expert's answer
2020-12-20T18:09:04-0500

We know that cosine law states that,

"a=\sqrt{b^2+c^2-2bc\cos\theta}"

and we have,

"b=12m"

"c=20m"

"\theta=45\degree"

"a=\sqrt{12^2+20^2-2\times 12 \times 20 \times \cos45\degree}"

"a=\sqrt{144+400-480 \times \frac{\sqrt2}{2}}"

"a=\sqrt{544 - 240\sqrt2}"

"a\approx 14.3"